11. Consider the following histogram of IQ scores of 7th graders. How many total students were included in the study?: * IQs12. Compute the sample standard deviation for (4, 5, 5, 6, 7, 7, 9, 9, 10, 12).: * O a. 0.8222 Ob. 0.8889 Oc. 2.2926Od. 2.5473 Questions 13-14: The following table displays the number of animal tracks collected in a park. Opossum Deer Squirr
Answer #1

(11) The Iq Scores of 7th graders are given to us. Score Interval Frequency (f) 90-99 60 100 - 109 80 60 110 - 119 120-129 40(4+5+5+6+7+7+9+9+10+12) To To (74) 10 7.4 Thus X = T. 4 M3 (x; - x2 Sample Standard deviation, S = n-1 x x =-7.4 (x-x)? 4 -3.Thus the samople standard deviation, S.2.5473 Thus the answer is (d) 2.5473 (13) Table for the number of animal tracks CollecPC a Yaodonly selected is a squirrel) - P(Squirrel) no of Squirrels Total no. of animals 97 217 0.4470 Thus P(arandomly selec(14) we have to find the a box & Whiskey plot. lower quartile to construct The data is given to us, ist we will arrange the gHere the median is the middle of the set of Ordered number. 15 and 32 Here the median is between Then the median mean of 15 a